OpenStudy (anonymous):

X+3y=z ? Work it out please

OpenStudy (baldymcgee6):

\[y = \frac{ z }{ 3 } - \frac{ x }{ 3 }\]

OpenStudy (baldymcgee6):

\[y = \frac{ z-x }{ 3 }\]

OpenStudy (anonymous):

Thanks!(:

OpenStudy (baldymcgee6):

please leave a best response :)